OpenStudy (anonymous):

Call Wireless phone company charges $0.50 for the first call you make and $0.10 for each additional minute. The cost of a phone call can be expressed by the formula C = 0.50 + 0.10(m – 1), where C is the total cost in dollars and m is the number of minutes that you talk. If the bill for your phone call was $1.90, how long did you talk?

OpenStudy (anonymous):

acc to ques. 1.90 = 0.50 +0.10(m-1) so 1.90 -1.50 / 0.10 =14 hence 14 + m-1 so m is equal to 15

OpenStudy (anonymous):

oh sorry 14 = m - 1 hence m = 15

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

you're welcome !!

OpenStudy (anonymous):

is it15